### v1

# Investigating the interaction of linguistic and mathematical reasoning in language models using multilingual number puzzles

Abstract
Across languages, numeral systems vary widely in how they construct and combine numbers. While humans consistently learn to navigate this diversity, large language models (LLMs) struggle with linguistic-mathematical puzzles involving cross-linguistic numeral systems, which humans can learn to solve successfully. We investigate why this task is difficult for LLMs through a series of experiments that untangle the linguistic and mathematical aspects of numbers in language. Our experiments establish that models cannot consistently solve such problems unless the mathematical operations in the problems are explicitly marked using known symbols ($+$, $\times$, etc, as in “twenty + three”). In further ablation studies, we probe how individual parameters of numeral construction and combination affect performance. While humans use their linguistic understanding of numbers to make inferences about the implicit compositional structure of numerals, LLMs seem to lack this notion of implicit numeral structure. We conclude that the ability to flexibly infer compositional rules from implicit patterns in human-scale data remains an open challenge for current reasoning models.
[…]
Investigating the interaction of linguistic and mathematical reasoning in language models using multilingual number puzzles
Antara Raaghavi Bhattacharya1, 2, Isabel Papadimitriou3, Kathryn Davidson2, David Alvarez-Melis1,3 1Computer Science Department, Harvard University,  2Linguistics Department, Harvard University, 3Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University Correspondence: antarabhattacharya@college.harvard.edu
1 Introduction
[…]
Variable Operator Example
Explicitness Familiarity Type
Single character Implicit - - A B
Explicit Familiar Symbol A + B
Explicit Unfamiliar Symbol A $\alpha$ B
Explicit Unfamiliar Word A xebrut B
Multi-character Implicit - - gbaifi pagig
Explicit Familiar Symbol gbaifi + pagig
Table 1: A demonstration of the experimental conditions for our explicit operators experiment. We add explicit operators to our base implicit problems, using both the familiar symbols for addition/multiplication/subtraction, as well as unfamiliar symbols and words to symbolize the operation.
We investigate the capabilities of language models to solve puzzles about linguistic number systems, drawn from linguistics competitions (Linguistics Olympiads) where high-school students have to reason through data about unknown languages and explain the linguistic rules governing the data Derzhanski and Payne (2010). While language models approach human performance on several language-based benchmarks Hendrycks et al. (2020); Kojima et al. (2022); Beguš et al. (2023), and recent reasoning models deliberately optimized for logical and mathematical reasoning show remarkable performance improvements for many structured mathematical reasoning tasks Zhong et al. (2024); Jaech et al. (2024), LLMs perform extremely poorly at solving linguistic-mathematical puzzles about systems of numbers in different languages Derzhanski and Veneva (2018); Bean et al. (2024).
Why do language models fail to solve these problems at the intersection of language and math — what specifically causes this failure? And how much of this failure is due to the linguistic vs. the mathematical aspects of the problem?
We present a method to systematically isolate individual parameters of number construction and combination and investigate how they affect language model performance. We establish that most individual mathematical features (like base) do not hinder the ability of sufficiently advanced language models to solve such problems. However, unless the mathematical operations in a problem are made explicit through familiar symbols (+, ×, etc.), models cannot consistently solve the problem. This indicates that, at least within the domain of linguistic-mathematical problems, models cannot infer the compositional structure of numerals like humans can, or sufficiently abstract notions like operators. We discuss our findings in the broader context of human language, concluding that flexible, adaptive use of language across domains appears to remain challenging for LLMs.
Figure 1: Making operators explicit significantly improves performance. Results for explicit operator experiments, for the single-character variable case (For the results on multi-character variables, see Appendix B Figure 7). Making operators explicit shows performance improvement over the implicit condition, but this is only substantially and reliably the case when the operator is made explicit with a familiar symbol like “+”. Error bars = standard error of the mean. 10 problems, 5 iterations per problem.
2 Methods
Models. We used OpenAI o1-mini Jaech et al. (2024) and DeepSeek-R1-distill-Qwen-7B Guo et al. (2025) reasoning models to conduct our experiments, querying o1-mini via the API and running DeepSeek locally. We will publicly release all code and data used for our experiments.
Data. We obtained data for linguistics olympiad problems from two publicly available datasets: LingOly Bean et al. (2024) and Linguini Sánchez et al. (2024), filtering both datasets for problems tagged as “number systems”. After filtering, we had 15 problems from the LingOly and 8 problems from the Linguini dataset. Not every problem in the dataset could be standardized in the ways that our experiments required. The entire dataset was thus manually evaluated for suitable problems, and 10 problems were chosen for evaluation, all in distinct languages (see Appendix D). These problems spanned a range of difficulty from the first round of the UK Linguistics Olympiad to the International Linguistics Olympiad (most challenging).
3 Experiments
3.1 The effect of explicit operators in problems
Since so many of the mathematical operators in numeral structure are implicit (eg, in English we say ‘twenty three’ to mean ‘twenty + three’), our first experiment investigates how this implicit structure affects how models solve the problems. To do this, we standardize and convert the 10 existing linguistic number system problems to mathematical problems, and vary how explicit the operators are, as shown in Table 1.
First, we standardize all problems to control for model tokenization and task-external knowledge effects: we identify all meaningful morphemes, standardize all phonological changes, and replace them with dummy words as described in detail in Appendix A. This standardized version of each problem is what we call the implicit setting, since the mathematical operations are largely implicit, as they are in language. Taking these implicit problems as our baselines, we then make the operators explicit in three ways: 1) as the familiar mathematical operator symbols that perform the operation (eg, ‘+’ for addition), 2) as symbols that are unfamiliar for performing that operation, and 3) as whole words sampled from the tokenizer. A full example prompt with a puzzle in four variations is provided in Appendix B.
We present our results in Figure 1. In all cases, the presence of explicit operations with familiar symbols yields significant improvements over the default implicit condition (o1-mini performs at ceiling). In the multi-character setting (more linguistic), models perform better on average in the implicit condition than in the case with an explicit operator as an unfamiliar random word (vid. Figure 7). It is likely harder to differentiate between function words (operators) and number words (numerals) in such a setting — this finding is consistent with work that has shown human solvers also find a problem to be more difficult when the operator word is explicit but unfamiliar Derzhanski and Veneva (2018). Overall, our results demonstrate that it is difficult for models to reason about the abstract that linguistic quantities might contain operators, if the operators are not explicitly provided using familiar symbols.
3.2 Providing contextual information
Our first experiment showed that in the absence of problem-specific instructions, when given a linguistic-mathematical problem directly, LLMs struggle to solve it unless the operations are both explicit and familiar. This leaves open the question of whether providing additional problem-specific information would affect the model performance. We thus modulate the context of the problem in three different ways. We query the same four problem variants as described in Table 1, additionally providing the following contextual information:
Language: “Here is a puzzle based on numbers in the {language} language."
Base: “Here is a puzzle based on numbers in a language that uses a base-{n} numeral system."
Implicit operations: “Here is a puzzle based on numbers in a language. In this language, numbers may be constructed through implicit operations like addition (twenty-nine = 20 + 9) or multiplication (five hundred = 5 $\times$ 100)." [only for implicit condition]
Figure 2: Language and base information only helps in the implicit case. Effect of adding language or numeral base information, plotted as a difference from the baseline values in Figure 1 for o1-mini. In cases with explicit operators, conflating overtly mathematical and linguistic information appears to confuse the models.
We compare these to the baseline results from Section 3.1 for o1-mini, presenting our results in Figure 2. In cases other than the implicit operator condition, the model seems to recognize the problem as requiring a more mathematical kind of reasoning, so providing linguistic information seems to confuse the model and average performance is worse. However, in the implicit operator (A B) condition, model performance improves significantly, perhaps because the setting of the problem is less overtly mathematical. In Figure 3, we show that providing information about the implicit reasoning needed is not as significant a boost as activating knowledge about the specific language.
Figure 3: Extra information improves performance on implicit problems (A B). Information about implicitness is helpful, but not as much as more direct information like the problem language. Error bars = standard error of the mean. 5 iterations / problem.
3.3 Ablations: constructed minimal-pair problems
In order to ensure that it is the difference in operators (as opposed to other features of the numeral system) that explains the models’ inability to solve these problems, we performed an ablation study to test whether models could handle other aspects of numeral construction and combination. Our experiment is inspired by the notion of a linguistic minimal pair, a pair of linguistic items that differ in exactly one meaningful element. We construct minimal pairs of simple, synthetic number system problems, where every element is the same except for one specific parameter that differs between two paired problems. We tested five major parameters of numeral systems, as described in Table 2.
Figure 4: Example of full minimal pair template problem, for the Order parameter, where we varied whether digits are read left-to-right or right-to-left.
In all cases, GPT-4 and more advanced models could solve the template problems. It thus appears that most basic “building blocks" of number systems (e.g. the base of the system, the order of numerals, etc.) did not affect model performance in isolation, but the models consistently fail to solve number problems that involve constructing and combining complex numerals.
Parameter GPT-$3.5$ GPT-$4$ o1111-mini
Numeral system vs. not AB = fifty one | AB = big bird
Typed vs. glyph AB = fifty one | AB = 51
Order L $\rightarrow$ R vs. L $\leftarrow$ R AB = 51 | BA = 51
Additive vs. subtractive AB = 27 (20 + 7) | AB = 27 (30 - 3)
Base of the numeral system*
Table 2: Minimal pair results: GPT-$4$ and o1-mini solve all paradigms, GPT-$3.5$-turbo struggles with numeral base and combination. Further data on testing all bases 4-19 linked in Table 5.
4 Discussion and Conclusions
We study the entanglement between linguistic and numeric knowledge in language models, focusing on the ability of models to use mathematical reasoning in problems that display the implicit numerical structure in language. In the setting of these linguistic-mathematical puzzles, we show that the overtness and familiarity of operators affects the performance of language models, although many humans are able to understand how numeral systems work and hence solve the problems without needing specified operators. However, a broader study with different controls and parameter settings remains open for future work. Since all our evaluation was standardized and closed-form, we welcome research on open-ended evaluation of reasoning task responses. Current language models seem to display some level of emergent modular structure Teehan et al. (2022); Lepori et al. (2023) — perhaps linguistic and mathematical tasks activate separate circuits or subspaces in models, and understanding the ways in which reasoning fine-tuning and reinforcement learning interacts with linguistic pretraining is another promising avenue for future research. Investigating such questions enriches our understanding of both computational and human approaches to representing numbers in language. The ability to understand language and abstract rule-governed systems is a fascinating aspect of human intelligence, and we hope that our research provides some insight into the understanding of this remarkable human trait.
Limitations
We acknowledge the possibility that our results are explained by limitations in the training data and the small size of our dataset, as language models often equal human performance on benchmarks for which they have large quantities of similar-enough training data Achiam et al. (2023). Perhaps an LLM trained on a massive corpus of linguistic number system problems would be able to solve new, previously unseen number system problems. But the data today are far too limited for such an approach, and crucially, a human solver who is familiar with existing number system problems can generalize to unseen problems extremely well! Even a human solver who is unfamiliar with existing number system problems can in theory solve any problem they are provided just by logically reasoning. Importantly, we note that although this may not be true of the average human, when comparing the top end of humans with the top-performing current language models, it is clear that intuiting rules from human-scale data is still challenging for LLMs.
Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University.
Appendix A Randomization strategies for task-external knowledge and tokenization handling
In this section, we address the specific changes we make across linguistic number system problems to convert them into templates suitable for our dataset. In order to truly test whether the model is solving a problem, it should not be affected by factors external to the problem, such as flawed tokenization or the usage of memorized knowledge external to the provided task. [^1] Memorized knowledge would also help a human solver, but people are much less likely to know the number systems of different (particularly low-resource) languages. Although linguistics olympiad contestants might know more number systems than the average person, there are over 7,000 human languages, so the probability of knowing a specific system is low. Moreover, since LLM training corpora scrape large portions of the internet, the breadth of their memorized knowledge far exceeds that of an average human.
Our strategy to remediate this issue is thus: in the single-letter token setting, we separated all characters by whitespaces to ensure correct tokenization. In the multi-token setting, we identified all meaningful morphemes in the problems and standardized them to remove any phonological changes, such that every morpheme had exactly one surface representation. We separated every meaningful morpheme with whitespaces, and mapped each morpheme to a randomly generated multi-token “dummy word" for each iteration of each experiment. We created each of these “dummy words" by randomly sampling short tokens (length $\leq 3$) from the language models’ respective tokenizer vocabularies, and concatenating tokens together to create unfamiliar words.
For tokenizers which use schemes like byte-pair encoding, any input string will get mapped to some sequence of tokens that are present in the vocabulary, so there is no situation in which the model will see an unknown token. Since the dummy words themselves have no meaning, the model cannot directly draw on task-external linguistic information to solve the presented problems. For simplicity we restricted the random draw to those containing only romanized (Latin alphabet) characters. We also excluded tokens that contained any numeral symbols from 0-9, to ensure that the the mathematical correctness of the problems was not affected.
Appendix B Multi-character-variable results from Exp 1
Figure 5: Drehu (IOL 2010) problem
We provide an example of our four variations of the puzzle in Table 3. To query all four variants, we used the same prompt “Here is a puzzle. Can you solve it? Please output only the answer (in place of the ??) and nothing else!".
Explicit + familiar
(masaad $\times$ pagig) + masaad + opbob = 31 (masaad $\times$ pagig) + masaad + buylen = 26 $\vdots$ (ajssci $\times$ pagig) + (ajssci $\times$ kould) = 50 (innops $\times$ pagig) + innops + opbob = ??
Implicit
masaad pagig nge masaad opbob = 31 masaad pagig nge masaad buylen = 26 $\vdots$ ajssci pagig nge ajssci kould = 50 innops pagig nge innops opbob = ??
Explicit + unfamiliar (Greek)
(masaad $\beta$ pagig) $\alpha$ masaad $\alpha$ opbob = 31 (masaad $\beta$ pagig) $\alpha$ masaad $\alpha$ buylen = 26 $\vdots$ (ajssci $\beta$ pagig) $\alpha$ (ajssci $\beta$ kould) = 50 (innops $\beta$ pagig) $\alpha$ innops $\alpha$ opbob = ??
Explicit + unfamiliar (random)
(masaad hibcat pagig) xebrut masaad xebrut opbob = 31 (masaad hibcat pagig) xebrut masaad xebrut buylen = 26 $\vdots$ (ajssci hibcat pagig) xebrut (ajssci hibcat kould) = 50 (innops hibcat pagig) xebrut innops xebrut opbob = ??
Table 3: Example of four problem variants in the multi-character setting, corresponding to Drehu (IOL 2010) dataset problem in Figure 5.
Appendix C Base experiment
In order to understand whether sufficiently advanced language models would show performance that was invariant to changes in the base, we conducted a more fine-grained minimal pair experiment into the effect of numeral base on problem performance. Here, the solver would see the Hindu-Arabic numerals corresponding to the English base-10 representation of the numbers, because the problem was presented in English. But the unknown symbols corresponded to the numbers as expressed in a different base, as shown in Figure 6.
Figure 6: Setup for base experiment
We conducted two different versions of this experiment. First, we mapped the unknown symbols to the single-character whitespaced A, B, C, and D tokens, as in Figure 6. In the second version, each of the four unknown symbols (A, B, C, D) was instead represented by a corresponding random token drawn from the tokenizer vocabulary, to ensure that the context of the specific tokens A, B, C, and D was not influencing our results.
We tested four increasingly sophisticated GPT models (GPT-$3.5$-turbo, GPT-$4$, GPT-$4$o, and o1111-mini) on both versions of the experiment and provide results in Table 5. GPT-4o and o1-mini solved all problems in both conditions, displaying performance that was robust to the base of the problem.
Appendix D Table of languages
The 10 problems that we used for our analyses. The problems range in difficulty from the first and second rounds of the UK Linguistics Olympiad (UKLO R1 and R2) to the International Linguistics Olympiad, which typically has the most challenging problems.
Language ISO code Base Level
Drehu dhv 20 IOL
Georgian kat 20 UKLO R1
Gumatj gnn 5 UKLO R1
Ndom nqm 6 IOL
Ngkolmpu kcd 6 UKLO R1
Northern Pame pmq 8 UKLO R1
Umbu-Ungu ubu 24 IOL
Waorani auc 5 UKLO R1
Yoruba yor 20 UKLO R2
Yup’ik esu 20 UKLO R2
Table 4: Languages and problem features in final dataset (after removing/standardizing phenomena)
Base GPT-$3.5$-turbo GPT-$4$ GPT-$4$o o1111-mini
ABCD Random ABCD Random ABCD Random ABCD Random
4
5
6
7
8
9
11
12
13
14
15
16
17
18
19
Table 5: Base experiment results: GPT-$4$o / o1111-mini solve every problem, regardless of randomization
Figure 7: Both o1-mini and DeepSeek struggle with the explicit-unfamiliar condition (o1 shows negative improvement, DeepSeek shows 0%) in the multi-character setting. Error bars = standard error of the mean. 5 iterations / problem tested for 10 problems.
Figure 8: Results per language, (a) single-character (b) multi-character: performance varies significantly by problem and operator type.